OpenStudy (anonymous):

I don't want the answer here so much as an explanation as to instructions explaining how to get there : the population of a colony of bacteria grows exponentially according to the function below (included in a minute), where t is the time in hours. How long until the population is 1000?

OpenStudy (anonymous):

\[b(t) = 12e^.2t\] (the t is part of the exponent but the formatting won't let me do that)

OpenStudy (anonymous):

so to get there, substitute b(t) with 100. so it looks like 1000=12e to the 2nd power multiplied by the t as a power \[1000=12e ^{2t}\] then you divide both sides by 12\[83.333333=e ^{2t}\] e is just a constant. which is 2.71828 so you get \[83.333333=2.71828^{2t}\] which means you need to square root both sides. \[\sqrt{83.333333}=2.71828^{t}\] which equals to \[9.1287092734953499558532073114273\] but we will round to 9.1287 \[9.1287=2.71828^{t}\] then you use the log formula-> \[\log_{2.71828}9.1287=t\] which means \[\log_{91287} / \frac{ \log _{10}9.1287 }{ \log _{10} 2.71828 }\] and the answer is 2.2114247842569241065266455869202 hours. or 2.2 hours

OpenStudy (anonymous):

Shouldn't this be posted in maths

OpenStudy (anonymous):

it's a written explanation, so it DOES count as writing